OpenStudy (anonymous):

Suppose we are given two functions, say r(x) and s(x). a) Is it true that r(x) + s(x) = s(x) + r(x) for all values of x? Why or why not? b) Is it true that r(x)

OpenStudy (anonymous):

i lost you.

OpenStudy (anonymous):

Addition and multiplication are both commutative (meaning you can switch the order of them)... It's not always true that r(0) = 0. if r(x) = x +1, then r(0) = 0+1 = 1

OpenStudy (anonymous):

That was a old problem i was working on. What happen to the last one?

OpenStudy (anonymous):

i was lagging...probably because of the size of the post :(

OpenStudy (anonymous):

we can continue here though...

OpenStudy (anonymous):

Oh. You have been so helpful!

OpenStudy (anonymous):

so let's recap...you need to find g(a) - g(a-b) + f(a)...

OpenStudy (anonymous):

where g(a) = 2-3x and f(x) = 3x+2-(4/x), right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and we have established that...

OpenStudy (anonymous):

g(a) = 2-3a and g(a-b) = 2-3a+3b, right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so...what is f(a)?

OpenStudy (anonymous):

and don't let the fraction scare you... ;)

OpenStudy (anonymous):

f(a)=3a+2-4/a?

OpenStudy (anonymous):

exactly!

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

now...let's put them together, distribute any negatives, and simplify.

OpenStudy (anonymous):

Ok so how do i write this out the right way?

OpenStudy (anonymous):

\[g(a)-g(a-b)+f(a)=(2-3x)-(2-3a+3b)+(3a+2-\frac{4}{a})\]

OpenStudy (anonymous):

distribute the negative in the middle and combine like terms...

OpenStudy (anonymous):

ack...change the x to an a....

OpenStudy (anonymous):

uhh. yikes. ok.

OpenStudy (anonymous):

i dont know.

OpenStudy (anonymous):

you should get\[2-3a-2+3a-3b+3a+2-\frac{4}{a}\]after changing the signs of the middle grouping...

OpenStudy (anonymous):

thats the answer?

OpenStudy (anonymous):

eanotice that each set of parenteses is just what we found to be g(a), g(a-b) and f(a)...

OpenStudy (anonymous):

no...now you need to combine like terms...combine things with an a together and combine things with a b together...

OpenStudy (anonymous):

anything that is just a number can combine as well..

OpenStudy (anonymous):

for instance, the first and second 2 cancel each other, right?

OpenStudy (anonymous):

-3b+3a+2-4/a?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

and tyou say you aren't good at math! i think you lied to me :)

OpenStudy (anonymous):

haha the fact that they canceled out was worryin me

OpenStudy (anonymous):

so now you got it, right? :)

OpenStudy (anonymous):

yes. wow. that helped so much.

OpenStudy (anonymous):

then it's time for me to depart...if you see me on and have more question, let me know.

OpenStudy (anonymous):

Thank you!! I will!!

OpenStudy (anonymous):

my pleasure.